OpenStudy (annipuppi):

what would be the sin, cos, and tan ratios of a triangle where the opposite side is 2.5, the adjacent side is 16, and the hypotenuse and 16.5?

OpenStudy (anonymous):

Remember SohCahToa? \[\sin \theta=\frac{ opp }{ hyp }\] \[\cos \theta=\frac{ adj }{ hyp }\] \[\tan \theta=\frac{ opp }{ adj }\]